c) A loan of £100 nominal is issued with interest payments (coupon) payable at 8% p.a. monthly in arrears for 7 years, with the amount of the loan repayable at the end of the term. For an investor who is liable to income tax at a rate of 22%, but not capital gains tax, with a required redemption yield of 6%pa., calculate the price the investor should pay for the loan. [3 marks]
Answer #1

Answer:

Given

Price of Loan Amount at redemption A=100

Coupon rate= 8%

Income tax T=22%

Coupon =100**8%*(1-22%)=6.24 per annum

Coupon per month C=6.24/12=0.52

Redemption yield=6 % p.a

Redemption yield per month r=6%/12=0.5%

Periods n=12*7=84 months

So Price of Loan P=C*{1-(1+r)^-n}/r + A/(1+r)^n

P=0.52*(1-(1+0.5%)^-84)/0.5% + 100/(1+0.5%)^84

P=101.37